The first angle of a triangle is 15 degrees more than the second triangle. The third angle of a triangle is three times the second angle. Use x to represent the unknown Value.
? The first angle
? the second angle
? the third angle
Answer #1
let x=second, so x+15=first and 3x=third x+(x+15)+(3x)=180
answered by: jockalyn quswbey
Answer #2
The first angle of a triangle is 15 degrees more than the second triangle. The third angle of a triangle is three times the second angle. Use x to represent the unknown Value let the second angle be x the first angle is then x+15 the third angle is 3 times the second so it is 3x all these add to 180 degrees (angle sum of triangle) x+(x+15)+3x=180 5x+15=180 minus 15 from both sides of the equals sign 5x+15-15=180-15 5x=165 divide both sides by 5 x=165/5 x=33 so the second angle is 33 the first is 33+15=48 the third is 3 times 33=99
answered by: Devan